OpenStudy (anonymous):

Find the foci and ellipse with the following equation: x^2/169 + y^2/25 = 1

OpenStudy (anonymous):

the form for this is \[\frac{x^2}{13^2}+\frac{y^2}{5^2}=1\]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

because 13 is larger than 5 you know it looks like this 

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the center is \((0,0)\) so to find the foci you need to go to the left and right from the origin (rather than up and down)

OpenStudy (anonymous):

\(a=13,b=5,c=\sqrt{13^2-5^2}=12\) i.e. you have a 5 - 12 - 13 right triangle

OpenStudy (anonymous):

Alright, got it

OpenStudy (anonymous):

therefore foci are 12 units to the left and right of the origin i.e. \((-12,0)\) and \((12,0)\)

OpenStudy (anonymous):

I got 0,-12 but I see my mistake